1.50 days to s,=

0.0600 mi to m=

1.89 $/gal to dollars per liter=

0.580 in/ms to km/hr=

25.00 gal/min to L/s=

0.05000 ft3 to cm3=

Express your answers using three significant figures.
Answer #1

1.) 1.5×24×60×60 = 129600 second

2.) 96.561 meters

3.) 1×1.89/3.785×1 = 0.499dollar per litre

4.) 0.58×91.44= 53.035km/hr

5.) 1.577L/s

6.) 0.05×28316.847 = 1415.842 cm3